OpenStudy (anonymous):

Have a question from an old textbook I just can't figure out. "Snail Races: As part of their preschool science project, Mrs. Joy Pribble's class is holding snail races. The first snail, Zippy, is known to move at a rate of 5 inches per hour. The second snail, Lightning, is known to move at 4.5 inches per hour. If the snails race along a straight path, and if Zippy finishes the race 0.25 hour before Lightning, a) determine the time it takes Lightning to run the race. b) determine the time it takes Zippy to run the race. c) what is the distance covered by both snails?"

OpenStudy (mathstudent55):

d = st Let t = shorter time (Zippy's time). Then Lightning's time is t + 0.25 Zippy: d = 5t Lightning: d = 4.5(t + 0.25) They covered the same distance, so d for both is the same. Then 5t = 4.5(t + 0.25)

OpenStudy (perl):

*

OpenStudy (anonymous):

Wow, I was making it whole lot more complicated than it was. I'm just rubbish at math, hopefully I can improve. Thank you sir.